OpenStudy (anonymous):

Select all of the answers that apply. Which of the following equations are equivalent to -2m - 5m - 8 = 3 + (-7) + m? -15m = -4m -7m - 8 = m - 4 -3m - 8 = 4 - m m - 4 = -7m - 8 -8 - 7m = -4 + m -8 - 3m = 4 - m

OpenStudy (anonymous):

-2m - 5m - 8 = 3 + (-7) + m -7m-8=-4+m -7m-8=m-4 second one...

OpenStudy (anonymous):

just the second one?

OpenStudy (anonymous):

and the fifth one .....after arranging it...

OpenStudy (anonymous):

k